OpenStudy (anonymous):

solve by elimination method 2a+4b=12 -2a+b=18

OpenStudy (anonymous):

What do you think the first step should be?

OpenStudy (uri):

@Avon just add both of the equations 2a+4b=12 -2a+b=18 ---------- What do you get?

OpenStudy (anonymous):

gypsy&@uri 4b=6

OpenStudy (anonymous):

@gypsy1274

OpenStudy (uri):

How? +4+1a=5a and 12+18=30

OpenStudy (uri):

I mean 5b*

OpenStudy (anonymous):

Break it into terms. \(2a-2a\) in the first column

OpenStudy (anonymous):

\(4b+b\) in the second column.

OpenStudy (anonymous):

i meant 5b=30

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Now, solve for b.

OpenStudy (uri):

2a+4b=12 -2a+b=18 ---------- 5b=30 Divide 30 with 5 b=30/5 b=6.

OpenStudy (uri):

Now you have the value of b now solve for a Put the value of b in eq no(1) 2a+4b=12 2a+4(6)=12 Solve it.

OpenStudy (anonymous):

now im lost but would it b 2a+24=12 @uri

OpenStudy (uri):

2a+24=12 Subract 24 both sides now..

OpenStudy (anonymous):

@uri 2a=12

OpenStudy (uri):

2a=-12 Now divide -12 with 2

OpenStudy (anonymous):

6

OpenStudy (uri):

-6 ans. so a=-6 and b is 6.

OpenStudy (anonymous):

thanks @uri